OpenStudy (anonymous):

Solve the equation 25x-4+17Squared+25y (3,19)

OpenStudy (anonymous):

What does the first equation equal?

OpenStudy (anonymous):

well the factor on 24x-4+17^2+25y= 5(5x+57+5y)

OpenStudy (anonymous):

can you simplify it?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

25x+25y+285

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

I don't need what x and y equal they are given I need what the equation equals if you plug them in

OpenStudy (anonymous):

so it would be 835 total

OpenStudy (anonymous):

75+475+285=835